OpenStudy (anonymous):

write an equation for the following, slope=2/3: (0,3)

OpenStudy (anonymous):

\[y=\frac{2}{3}x+3\]

OpenStudy (anonymous):

y=mx+b has slope m and y-int b or (0,b)

OpenStudy (anonymous):

good?

OpenStudy (anonymous):

Thank you, so for slope -4: (0,5), the equation would be y=-4x+5 ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Thank you, would you happen to know if y=2/3x is a linear relationship?

OpenStudy (anonymous):

yes; if you are able to write the eqation in the form y=mx+b or y=b or ax+by=c or x=a then you have a linear relationship; a, b,c are real numbers

OpenStudy (anonymous):

they all produce lines; lines are "linear"

OpenStudy (anonymous):

Yep, that makes sense, thank you..

OpenStudy (anonymous):

ur welcome :})